OpenStudy (anonymous):

Let f(x)= –5x + 3 and g(x) = 6x – 2. Find f•g and its domain. 1)–30x^2 + 28x – 6; all real numbers except x=1/3 2)–15x^2 – 8x – 12; all real numbers except x= 3/5 3)–30x^2 + 28x – 6; all real numbers 4)–15x^2 – 8x – 12; all real numbers Please explain how you got it.

OpenStudy (anonymous):

you must solve ( –5x + 3)*(6x – 2)

OpenStudy (anonymous):

Do i multiply -5x by 6x and 3 by 2

OpenStudy (anonymous):

(a+b)(c+d)=a(c+d)+b(c+d)=ac+ad+bc+bd

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

plz wait

OpenStudy (anonymous):

yes that's true

OpenStudy (anonymous):

ok thanks